OpenStudy (anonymous):

The length of a rectangle plus its width is 24 cm. The area is 143 square cm. What are the length and width of the rectangle?

OpenStudy (anonymous):

Let length be L. Length + Width = 24 L + Width = 24 Width = 24 - L Area of rectangle = Length x Width = 143 143 = (L)(24 - L) 143 = 24L - L^2 Factorize / do quadratic equation solver to find L. As you would obtain two values of L, but L > 0 to eliminate one of those values. Hope this helps.

OpenStudy (anonymous):

thanks :)